OpenStudy (anonymous):

can anyone tell me what the points (40,400) (50,475) in standard form are?

OpenStudy (anonymous):

I know that the slope is 15/2 .... and the equation is y-y1 = m(x-x1) I just want to make sure I am correct before I submit.... y-400 = 15/2(x-40) y-400 = 15/2x - 300 y-400 - 15/2x = 300 y + 15/2 x = 100 is this correct in standard form? y+15/2x = 100

OpenStudy (anonymous):

I believe standard form is when the thing equals 0.

OpenStudy (anonymous):

so y+15/2 x - 100=0

OpenStudy (anonymous):

But yep looks right :) .

OpenStudy (anonymous):

are you sure? @Dido525

OpenStudy (anonymous):

Nvm, your answer is correct :) . Good job! :) .

OpenStudy (anonymous):

check your messages! thank you (: @Dido525